Calculate 1. 25 *10-1* (-4.5 * 10-1+ 12.5 * 10-2) by hand, assuming each of the values is stored in a single precision format. Show all your steps,similar way to the examples you took in class,and write your answers in both the single-precision floating-point format and in decimal.
Answer #1
1.25 *10-1* (-4.5 * 10-1+ 12.5 * 10-2)
= 1.25 *10-1* (-45-1+ 12.5 * 10-2)
= 1.25 *10-1* (-45-1+ 125-2)
= 1.25 *10-1* (-46+ 125-2)
= 1.25 *10-1* (79-2)
= 1.25 *10-1* (77)
= 12.5-1* (77)
= 12.5-77
= -64.5